OpenStudy (anonymous):

What is 3 √250 expressed in simplest radical form?

mathslover (mathslover):

\[3\sqrt{25*2*3}\] \[3*5\sqrt{6}\] \[15\sqrt{6}\]

OpenStudy (unklerhaukus):

@mathslover \[25\times2\times3\neq25\times2\times5=25\times10=250\]

mathslover (mathslover):

opps sorry @UnkleRhaukus ... mistake there so it will be \[15\sqrt{10}\]

OpenStudy (unklerhaukus):

yeah

OpenStudy (anonymous):

\[15\sqrt{10}\]